OpenStudy (anonymous):

Which of the following is the mathematical representation of Gay-Lussac's law? (Points : 3) P1V1 = V2/T2 P1/T1 = P2/T2 V1/T1 = V2/T2 P1V1 = P2V2

OpenStudy (anonymous):

@Somy I chose B

OpenStudy (somy):

you are right :3

OpenStudy (anonymous):

Thank you! ^_^

OpenStudy (anonymous):

A student increases the temperature of a 300 cm3 balloon from 30°C to 60°C. What will the new volume of the balloon be? (Remember to be careful with the units.) (Points : 3) 273 cm3 300 cm3 600 cm3 329 cm3

OpenStudy (somy):

which formula would you use for this? you have volume and temperature

OpenStudy (anonymous):

i have no clue

OpenStudy (somy):

okay remember this formula also : Charles Law \[\sf \frac{ V_1 }{ T_1 } = \frac{ V_2 }{ T_2 }\]

OpenStudy (somy):

you are given V1 T1 T2 you have to find V2 units for this formula are: V = m3 T = Kelvin

OpenStudy (anonymous):

so whats v1 t` and t2?

OpenStudy (somy):

V 1 is initial volume T1 is initial temperature

OpenStudy (somy):

V2 is final volume and T2 is final temperature

OpenStudy (anonymous):

Okay so I would have to do the fishtail division we learned in third frade for this?

OpenStudy (anonymous):

grade**

OpenStudy (somy):

I always choose the safe zone when it comes to units which is why i told you to use the SI unit system so first convert the units to ones i told you and then calculate using that formula & yes do this 3rd grade calculation AFTER converting to appropriate units

OpenStudy (somy):

your final answer for volume will be in m3 so convert it to cm3 to match one of the options

OpenStudy (anonymous):

i am so so confused

OpenStudy (anonymous):

is it C?

OpenStudy (somy):

Confused? lol u said grade 3 calculation :P okay look T is given to you in Celsius and to convert it to Kelvin all you need to do is add 273 to Celsius value

OpenStudy (somy):

1cm3 = 1.0 × 10-6 m3 300cm3 = X? m3 X= 300 × ( 1.0 × 10-6) so find it

OpenStudy (somy):

once you find it \[\sf \frac{ V_1 }{ T_1 } = \frac{ V_2 }{ T_2 }\] you need V2 so rearrange the formula \[V_2 = \frac{ V_1 \times T_2 }{ T_1 }\]

OpenStudy (somy):

try solving now

OpenStudy (anonymous):

ITS STILL CONFUSING. SO VERY CONFUSING. I dont get it. =.=

OpenStudy (somy):

okay, which part you dont get?

OpenStudy (anonymous):

From the very beginning to the very end

OpenStudy (somy):

do you get the formula?

OpenStudy (anonymous):

What I get is I have to plug things in to get the answer. thats it

OpenStudy (somy):

so you dont get the units conversion

OpenStudy (somy):

okay i'll make it easy for you since you dont get units forget about unit of volume leave it as cm3 only change the temp 30°C +237 = 303 K = T1 60°C +273 = 333 K = T2 V1 = 300 cm3 use the formula do grade 3 calculation \[\sf V_2 = \frac{ V_1 \times T_2 }{ T_1 }\] \[\sf V_2 = \frac{ 300 \times 333 }{ 303 }\]

OpenStudy (somy):

does it make sense now?

OpenStudy (anonymous):

sorry the internet blew out

OpenStudy (somy):

its okay, more importantly, did you get it?